# Conditional Denoising Diffusion Probabilistic Models for Data Reconstruction Enhancement in Wireless Communications

### v2

@@ IV-B Reconstruction Performance Metrics
Figure 6: Reconstruction performance in terms of peak SNR (PSNR) for different training strategies.
Fig. 6 shows the reconstruction performance in terms of peak signal-to-noise ratio (PSNR) metric for our proposed scheme. For case I to III, we consider $(\kappa^{t},\kappa^{r})=(0.01,0.01)$, $(\kappa^{t},\kappa^{r})=(0.05,0.1)$, and $(\kappa^{t},\kappa^{r})=(0.05,0.15)$, respectively. In this figure, we study the inference performance of the employed conditional DDPM when trained with two different training strategies: varying training SNR, or fixed training SNR. As mentioned earlier, the training samples are obtained by transmitting the training set of MNIST over the Sionna’s simulator (treated as the samples of ${\bm{x}}_{0}$) and obtain the decoded signals at the receiver’s output (treated as the samples of $\widehat{\bm{x}}_{0}$). This can be done with either fixing the communication scenario to a determined SNR, or simulating the system with different transmit SNRs. For the varying scenario, we set $-15\leq\Gamma^{\sf train}\leq 15$ dB by trail and error over different ranges of SNRs. Notably, when the conditional diffusion model is trained using training samples that are obtained over a wide range of communication SNRs, this makes the system more robust to the unseen scenarios in the inference phase. Therefore, as can be seen from the figure, higher PSNR values can be achieved in this case, while the performance of the fixed training SNR scenario saturates too early. We further note that this saturation is also due to another fact as well, which is known as “ceiling phenomena” in the literature of hardware-impaired communications [21]: By increasing the SNR, the effect of channel noise becomes negligible compared to the effect of hardware impairments, hence the performance metrics become saturated. As can be seen from the figure, the PSNR increases with the increase in SNR, due to having less noise communication, which makes it easier for the employed diffusion to denoise and reconstruct when facing with a less degraded signal. In addition, lower impairment levels indicate higher qualities for the transceiver hardware which results in less mismatches between the transmitted signals and the decoded ones, thus improving the reconstruction performance. Generally speaking, the proposed scheme shows a near-invariant performance with respect to the hardware impairments (as we can see the slight change in PSNR with respect to the three scenarios). This is further studied in the subsequent figures as well.
Figure 7: Reconstruction performance in terms of peak SNR (PSNR): Comparison between generative (DDPM-based) learning approach vs. discriminative (DNN-based) learning approach.
@@
Fig. 7 evaluates the reconstruction performance of our proposed approach when it is employed in conjunction with the Sionna’s neural receiver [^4] https://nvlabs.github.io/sionna/examples/Sionna_tutorial_part4.html, compared to the non-DDPM approach which simply employs the Sionna’s neural receiver without any diffusion enhancement. For the DNN-based receiver, we consider the network architecture proposed in [18], trained for 5000 iterations with learning rate $0.01$. Since we have employed our diffusion model at the receiver, we only exploit the receiver DNN of [18], and fine-tune it for benchmarking, so that we can have a fair comparison. For the DNN benchmark, three hidden linear layers with $64$ neurons and rectified linear unit (ReLU) activation functions are considered. Based on our trail and errors, increasing the depth of the conventional DNN, or the number of hidden neurons does not result in any significant improvement in its performance. The figure highlights about $10$ dB improvement in low-SNR regimes, as well as more than $5$ dB improvement in higher SNRs, when using our generative model. This is due to the fact that our DDPM-based scheme takes into account the entire underlying distribution of the ground-truth signals that are supposed to be sent over the wireless channel. However, the conventional DNNs neglect the underlying “content” of the information signals, and simply care about the “bit-wise” transmission, using discriminative loss functions, such as binary cross-entropy, rather than a content-wise” approach.
@@
Fig. 9 shows the reconstruction performance between the original images and the reconstructed ones for two different impairment levels. In this experiment, we aim to show that the functionality of the proposed conditional diffusion model can be exploited to compensate for the underlying errors within the noisy reconstructed images obtained at the receiver, instead of employing channel codes. As can be seen from the figure, by employing our conditional DDPM, the quality of the information signals (images in our scenario) can be improved more than the case where error correction codes are employed. For instance, by employing our proposed scheme, more than $10$ dB improvement in reconstruction performance can be achieved at $5$ dB SNR and impairment level $0.1$. This can be explained due to the fact that channel codes do not take into account the underlying structure and the contents of the information signals that are transmitted over the channel. However, our proposed scheme essentially considers the content of the transmitted signals when carrying out the denoising and reconstruction. Also note that error correction codes require sending redundancy bits which decrease the data rate, while our diffusion-based scheme has achieved better performance without the need for transmitting such redundancy bits. Therefore, the figure indicates that our scheme can improve the reliability of the communication systems, while saving the resources for only transmitting information bits. Obviously, this performance improvement comes with a trade-off in the sense that the DDPM model requires higher computation resources than the error correction codes. Nevertheless, one should note that the computations related to error correction codes are implemented in both of the communication sides, i.e., the transmitter (the encoder part), and the receiver (the decoder part), while the proposed diffusion is only employed and run at the receiver side. We note that studying the mathematical models and expressions for this trade-off is left for our future works.
@@
Fig. 10 studies the effect of different HWI levels on the reconstruction performance of our scheme over Rayleigh fading channels. For this figure, we set $\kappa^{t}=0.05$ and vary the impairment level of the receiver, $\kappa^{r}$, over the typical ranges specified in Ref. [21]. The reconstruction results are obtained in terms of MSE metric over $20$ realizations of the system. The figure highlights an important characteristic of our proposed scheme. Our DDPM-based communication system is resilient against hardware distortions, since the reconstruction performance does not change with the increase in the impairment level. Notably, our scheme showcases a near-invariant reconstruction performance with respect to noise and impairment levels. This is achieved due to the so-called “variance scheduling” of diffusion framework in (3), which allows the system to become robust against a wide range of distortions caused by channel and hardware impairments.
V Conclusions & Future Works
@@
In this paper, we have proposed conditional DDPMs to enhance the data reconstruction in wireless communication schemes. The key idea is to leverage the generative prior of diffusion models to learn a noisy-to-clean transformation of the information signal, conditioned on the degraded version of the receiver data. The proposed scheme is applicable to scenarios in which a prior knowledge of the information signals is available, e.g., in multimedia transmission. Our numerical results on MNIST dataset have shown that instead of employing complicated channel codes that reduce the information rate, conditional diffusions can be employed that improve reliable data transmission, especially under extreme channel conditions due to low signal-to-noise ratio (SNR), or hardware-impaired communications. Our evaluations have been carried out using the NVIDIA Sionna simulator, where it has been shown that more than $10$ dB improvement as well as $2$ times higher data rates can be achieved under low SNR regimes with our proposed scheme, compared to a conventional error correction-based data transmission and reconstruction. Moreover, we highlight the outperformance of the proposed generative learning approach compared to the conventional contrastive learning models based on DNNs. One important direction to be studied in the future works is to analytically study the complexity of generative AI-based approach versus the normal heavily channel coded approach.

### v3

@@ IV-D Quantitative Comparisons: Reconstruction Performance
Figure 7: Reconstruction performance in terms of PSNR for DDPM vs. VAE.
Fig. 7 shows the reconstruction performance in terms of PSNR metric for our proposed conditional DDPM-based approach, compared to the VAE-based reconstruction baseline which was introduced in Section IV-B. For case I to III, we consider $(\kappa^{t},\kappa^{r})=(0.01,0.01)$, $(\kappa^{t},\kappa^{r})=(0.05,0.1)$, and $(\kappa^{t},\kappa^{r})=(0.05,0.15)$, respectively. In the figure, VAE-Dim shows the latent space dimension of the implemented VAE, where higher latent dimensions can offer a slightly better reconstruction quality. As can be seen from the figure, our proposed DDPM outperforms the VAE baseline with more than $40\%$ improvement in PSNR metric. This is due to the high-quality sampling characteristic of diffusion models, while VAEs lack this important generative feature [28]. To elaborate, in contrast to VAEs, diffusion models maintain high sample quality and strong mode coverage, because of their unique feature for fine-grained denoising which is realized via the reverse diffusion process. This makes them suitable for the applications that interact with end-users and require high quality data reconstruction and generation. We further note that the saturation observed in the PSNR trends over SNR is due to the so-called “ceiling phenomena” in the literature of hardware-impaired communications [21]. To elaborate, by increasing the SNR, the effect of channel noise becomes negligible compared to the effect of residual impairments within the RF chain of communication systems, hence the performance metrics become saturated [21]. As can be seen from the figure, the PSNR increases with the increase in SNR, due to having less noise communication, which makes it easier for the employed diffusion to denoise and reconstruct when facing with a less degraded signal. In addition, lower impairment levels indicate higher qualities for the transceiver hardware which results in less mismatches between the transmitted signals and the decoded ones, thus improving the reconstruction performance. Generally speaking, the proposed scheme shows a near-invariant performance with respect to the hardware impairments (as we can see the slight change in PSNR with respect to the three scenarios). This is further studied in Fig. 11.
Figure 8: Reconstruction performance in terms of PSNR: Comparison between generative (DDPM-based) learning approach vs. discriminative (DNN-based) learning approach.
@@
Fig. 8 evaluates the reconstruction performance of our proposed approach when it is employed in conjunction with the Sionna’s neural receiver [^7] https://nvlabs.github.io/sionna/examples/Sionna_tutorial_part4.html, compared to the non-DDPM approach which simply employs the Sionna’s neural receiver without any diffusion enhancement. The figure highlights about $10$ dB improvement in low-SNR regimes, as well as more than $5$ dB improvement in higher SNRs, when using our generative model. This is due to the fact that our DDPM-based scheme takes into account the entire underlying distribution of the ground-truth signals that are supposed to be sent over the wireless channel. However, the conventional DNNs neglect the underlying “content” of the information signals, and simply care about the “bit-wise” transmission, using discriminative loss functions, such as binary cross-entropy, rather than a content-wise” approach.
@@
Fig. 10 shows the reconstruction performance between the original images and the reconstructed ones for two different impairment levels. In this experiment, we aim to show that the functionality of the proposed CDiff can be exploited to compensate for the underlying errors within the noisy reconstructed images obtained at the receiver, instead of employing channel codes. As can be seen from the figure, by employing our conditional DDPM, the quality of the information signals (images in our scenario) can be improved more than the case where error correction codes are employed. For instance, by employing our proposed scheme, more than $10$ dB improvement in reconstruction performance can be achieved at $5$ dB SNR and impairment level $0.1$. This can be explained due to the fact that channel codes do not take into account the underlying structure and the contents of the information signals that are transmitted over the channel. However, our proposed scheme essentially considers the content of the transmitted signals when carrying out the denoising and reconstruction. Also note that error correction codes require sending redundancy bits which decrease the data rate, while our diffusion-based scheme has achieved better performance without the need for transmitting such redundancy bits. Therefore, the figure indicates that our scheme can improve the reliability of the communication systems, while saving the resources for only transmitting information bits. Obviously, this performance improvement comes with a trade-off in the sense that the DDPM model requires higher computation resources than the error correction codes. Nevertheless, one should note that the computations related to error correction codes are implemented in both of the communication sides, i.e., the transmitter (the encoder part), and the receiver (the decoder part), while the proposed diffusion is only employed and run at the receiver side.
Remark 3
We remark that the outperformance of our scheme compared to the traditional channel codes comes with a trade-off. DDPMs may fall short in sampling speed due to the stepwise denoising process. Although this appears to limit the wide adoption of diffusion models for some applications, there already exists an active line of research focusing on accelerating the generation process. Such computation-efficient variants of diffusion framework, such as [4], can help facilitate the applicability of diffusion models to a wide range of real-world wireless scenarios.
IV-E Complexity Analysis
We further address the potential computational overhead of implementing DDPMs to highlight the complexity-performance trade-off compared with the traditional error correction coding methods.
IV-E1 Traditional channel codes
According to the coding theory, the encoding of LDPC codes is typically performed using a generator matrix $\mathbf{G}$ that transforms the input information bits into codewords via multiplying the input stream by the 5G LDPC matrices. The complexity of encoding an LDPC code is proportional to the number of non-zero entries in the generator matrix, and is approximated by $\mathcal{O}(n)$, where $n$ is the length of the codeword. The complexity of LDPC decoding is more computation-intensive than encoding, because of the iterative algorithms of decoding. According to coding theory [29], the complexity of LDPC decoding depends on the number of non-zero entries in the parity-check matrix $\mathbf{H}$, which represents the edges in the bipartite graph used for message passing in the belief propagation (BP) algorithm [29, 30]. The BP algorithm iteratively updates the likelihoods (beliefs) for each bit, passing messages between the variable nodes (representing bits) and check nodes (representing parity checks) in a bipartite graph. Then the complexity per iteration is proportional to the number of edges, $E$, in the graph. For an LDPC with the length of the codeword denoted by $n$, the complexity per iteration is approximately $\mathcal{O}(E)$, and since the number of edges $E$ is typically proportional to $n$ (due to the sparsity of $\mathbf{H}$), the computation complexity is $\mathcal{O}(n\cdot I)$, where the number of iterations $I$ for LDPC decoding typically depends on the target error-correction performance. Assuming $F$ floating-point operations (FLOPs) per message update on an edge (for calculating likelihoods and probabilities), the total FLOPs for the traditional error correction would be
$\text{FLOPs}^{\mathsf{LDPC}}=\mathcal{O}(I\cdot n\cdot F).$ (35)
IV-E2 DDPM’s complexity
The computational complexity of a U-Net can generally be analyzed in terms of the number of FLOPs for its convolutional layers within the contractive and expansive paths [26, 31]. Key parameters that influence the complexity are input size (considering the spatial dimensions of the input image as $H\times W$ for height and width), and depth of the U-Net architecture, $D$. This reflects the number of down-sampling and up-sampling stages in the U-Net. For a U-Net with depth $D$, the input size is halved at each down-sampling layer in the contracting path, and the number of channels doubles [26]. Hence, assuming $k\times k$ filter size and $C$ as the number of input channels, the FLOPs at each layer $i\in[D]$ of the contractive (down-sampling) path can be calculated as
$\text{FLOPs}^{\mathsf{UNet,ds}}_{i}=k^{2}\left(\frac{H}{2^{i}}\cdot\frac{W}{2^{i}}\right)\cdot(2^{i-1}C)\cdot(2^{i}C).$ (36)
The reverse happens in the up-sampling (expansive path), and hence we have
$\text{FLOPs}^{\mathsf{UNet,us}}_{i}=k^{2}\left(\frac{H}{2^{D-i}}\cdot\frac{W}{2^{D-i}}\right)\cdot(2^{D-i}C)\cdot(2^{D-i-1}C).$ (37)
In addition to up/down-sampling, U-Net architecture also maintains the skip connections, the complexity of which is negligible with respect to that of the convolution layers. Hence, summing over all the down-sampling and up-sampling layers for a depth $D$ U-Net, the total complexity of the U-Net architecture can be approximated by [31]
$\text{FLOPs}^{\mathsf{U-Net}}\approx\mathcal{O}(H\cdot W\cdot C^{2}\cdot D).$ (38)
Recall that in the denoising and reconstruction process of our proposed CDiff method, the U-Net is applied at each of the $T$ diffusion steps to gradually denoise the data. At each step, the complexity is that of running the U-Net once. Therefore, the total complexity of the diffusion model is the complexity of the U-Net multiplied by the number of diffusion steps
$\text{FLOPs}^{\mathsf{DDPM}}\approx\mathcal{O}(T\cdot H\cdot W\cdot C^{2}\cdot D).$ (39)
This indicates that the complexity grows linearly with the depth $D$ of the network, and linearly with the spatial size of the input $H\times W\times C^{2}$.
IV-E3 Comparison
If we assume the number of denoising steps $T$ to be equivalent/counterpart to the number of decoding iterations, $I$, in traditional decoding, and the input spatial dimensions $H\times W\times C^{2}$ in the DDPM method as the equivalent counterpart to the size of the input data-stream to the LDPC error correction decoding (input codeword length), the computation complexity of the DDPM-based approach depends on how deep the neural architecture of the diffusion framework is. This is reflected as the depth $D$ of the network in (39). This shows that, similar to any other AI/ML models applied to wireless problems, the complexity overhead of diffusion-based models is also dominated by nothing except the underlying neural network architecture. This also highlights the importance of studying efficient implementations of diffusion architectures and quantization and pruning techniques for the wide adoption of such models in real-world wireless problems.
@@ IV-E3 Comparison
Finally, Fig. 11 studies the effect of different impairment levels on the reconstruction performance of our scheme over Rayleigh fading channels. For this figure, we set $\kappa^{t}=0.05$ and vary the impairment level of the receiver, $\kappa^{r}$, over the typical ranges specified in Ref. [21]. The reconstruction results are obtained in terms of MSE metric over $20$ realizations of the system. The figure highlights an important characteristic of our proposed scheme. Our DDPM-based communication system is robust against hardware distortions, since the reconstruction performance does not change with the increase in the impairment level. Notably, our scheme showcases a near-invariant reconstruction performance with respect to noise and impairment levels. This is achieved due to the so-called “variance scheduling” of diffusion framework in (3), which allows the system to become robust against a wide range of distortions caused by channel and hardware impairments.
IV-F Further analysis and discussions on model’s robustness
We would like to provide some further discussions and highlight how the proposed CDiff method offers a robust performance under varying conditions, demonstrating its reliability in practical scenarios. The fact that the model adapts to varying noise conditions and is effective in low-SNR regimes can be addressed by the compilation of the following reasons.
  • DDPMs maintain the variance scheduling in the forward diffusion mechanism, wherein the samples are diffused at each time-step $t\in[T]$ according to $q(\bm{x}_{t}|\bm{x}_{t-1})$ in (1). Invoking (3)–(6), one can model the forward diffusion according to a “diffusion SNR” $\gamma^{\mathsf{DDPM}}(t)=\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}.$ The mathematical relations between the diffusion SNR, and the communication SNR would be an interesting direction to study in the future. However, the important point here is that in practice, the receiver of a communication system might not have any prior knowledge regarding the levels of RF impairments and communication SNRs. The forward diffusion does not consider any form of pre-defined hardware impairment levels or communication SNRs either. This is exactly the point of forward diffusion in our scheme. We purposefully “diffuse” Gaussian noise which maintains the maximum entropy, thus maximum uncertainty from the information-theoretic perspective, and then the DDPM tries to make itself “robust” against a wide range of uncertainties. This intuitively reflects a robust performance over a wide range of scenarios as shown in the paper.
  • The concept of conditioning in our scheme offers an additional level of robustness for the system. We incorporated the noisy decoded data $\widehat{\bm{x}}$ into the diffusion framework to learn to remove the exact noise from $\widehat{\bm{x}}$ and recover the ground-truth $\bm{x}$. This way, the CDiff learns to estimate both the Gaussian noise $\epsilon$ and the residual errors ${\widehat{x}}-{x}_{0}$. Moreover, this further allowed us to consider different communication SNRs during the training as described in IV-A. Hence, this conditioning has helped us incorporate the conditions of a practical wireless system into account, making the system more robust to different conditions in the inference phase.
  • The notion of time embedding in the DDPM also reinforces the model robustness. Considering the variance scheduling $\beta_{t}$, each time-instance has a unique mapping to a specific noise variance. Hence, intuitively speaking, each time-step equivalently corresponds to a specific level of noise-plus-distortion in the batch of received signals. Incorporating the time-steps $t\in[T]$ into our neural network as explained in Section IV-A, makes the model “know” at which particular time-step is operating for each level of channel noise, making it robust against different SNR and/or impairment levels. This notion $t$ is also reflected in the loss function of our model in (35), so data samples with different levels of “noisiness” are processed at different time-steps during training.
V Conclusions
@@
In this paper, we have proposed conditional DDPMs to enhance the data reconstruction in wireless communication schemes. The key idea is to leverage the generative prior of diffusion models to learn a noisy-to-clean transformation of the information signal, conditioned on the degraded version of the receiver data. The proposed scheme is applicable to scenarios in which a prior knowledge of the information signals is available, e.g., in multimedia transmission. Our numerical results on MNIST dataset have shown that instead of employing complicated channel codes that reduce the information rate, CDiffs can be employed that improve reliable data transmission, especially under extreme channel conditions due to low SNR, or hardware-impaired communications. Our evaluations have been carried out using the NVIDIA Sionna simulator, where it has been shown that more than $10$ dB improvement as well as $2$ times higher data rates can be achieved under low SNR regimes with our proposed scheme, compared to a conventional error correction-based data transmission and reconstruction. Moreover, we highlight the outperformance of the proposed generative learning approach compared to the conventional contrastive learning models based on DNNs.